Session: May 1998 (session 2)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Torun (K. Borkowski at Wed Jun 10 10:10:06 METDST 1998) Success
No known problems.

Onsala (A. Polatidis at Wed Jun 10 10:57:19 METDST 1998) Failure
The telescope was stowed due to high winds (>13-15m/s).

Noto (C.Trigilio at Wed Jun 10 11:21:58 METDST 1998) Success

Noto (C.Trigilio at Wed Jun 10 11:22:03 METDST 1998) Success

Urumqi (Liu Xiang at Wed Jun 10 11:45:02 METDST 1998) Success
The Clock Offset is 1.7us. It is cloud.

Westerbork (Tony Foley at Wed Jun 10 11:53:41 METDST 1998) Minor failures
No measurements of weather info or maser-gps from start to 14:51UT ; formatter to gps offset is continually monitored. Very strong EMI from about 14:40 to 15:20UT, but mostly in IF2.

Jodrell (Alastair Gunn at Wed Jun 10 12:27:37 METDST 1998) Minor failures
Problems with telescope control at start, FS restarted many times. Telescope then off-line due to high winds. Thus no data taken until about 12:42 UT. Strong RFI as usual.

Calibration (Jean\-Francois Desmurs at Thu Jun 11 15:24:43 METDST 1998) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the may98 directory under the name el019.antab. Read the comments in the file to get further information.

Effelsberg (dgraham at Fri Jun 12 15:05:40 METDST 1998) Minor failures
1841 off srce for a while with antenna problems